Use the figures in the table for the calculation. What is the NNP?

Government Purchases $158 billion
Depreciation $28 billion
Consumption $427 billion
Investment $43 billion
Exports $87 billion
Imports $103 billion
Income receipts from rest of the world $7 billion
Income payments to rest of the world $5 billion
$612 billion
$858 billion
$586 billion

Which is a factor that drives economic growth?

Decreased unemployment.
Increased human capital.
Improved investment opportunities.

The Helper Company’s customer service representatives received new computers and training on how to utilize them best. In regards to measuring productivity, these new computers and updated technology knowledge have

increased capital deepening for the Helper company.
increased profit for the Helper company.
increased costs for the Helper company.
Answer #1

1. NNP= GNP-depreciation.

GNP includes GDP + Net factor income from abroad.

In other words GDP is C+I+G+(X-M). when we add net factor income to the GDP we get GNP ie C+I + G + (X-M) + NFIA. By deducting the depreciation from GNP we get the NNP.

Here GDP= Consumption=$427 billion+ Investment =$43 billion+ Government purchase=$158 billion+ Net export(X-M, 103-87)-16=612

GNP= GDP+NFIA( Factor income receipts from the rest of the world-income payments to the rest of the world) = 612+2($7billion-$5billion) =614

NNP = GNP- Depreciation i.e $614billion-$28 billion=$586 billion.

Answer: $586 billion.

2. The basic factors that promote economic growth are: Natural resources, Physical capital formation, increased human capital and technology.

The improvement in human capital formation improves the quality of labour force. The labours get more skills, ability and training. Thus a skilled and trained labour force promotes growth since the workers are more productive.

Answer: Increased human capital.

3. Capital formation is the net addition to the existing physical stock of capital. The purchase of new computers by the company increases the physical capital and the training improves the human capital.

Answer: Increased capital deepening for the Helper Company.